OpenStudy (anonymous.):

Algebra 1 Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6. y = 3x + 7 y = 3x - 11 y = -3x + 11 y = -3x - 12

OpenStudy (anonymous):

parallel lines and slope. what should the slope of a parallel line be ?

OpenStudy (anonymous.):

slope = 3x ?

OpenStudy (anonymous):

correct !

OpenStudy (anonymous):

now just see which one satisfies the point

OpenStudy (anonymous.):

:) !! y-2=3(x-3) ?

OpenStudy (anonymous):

y-y0 = m(x-x0) P = (3, -2) y-[-2] = 3(x-[3]) y+2 = 3(x-3)

OpenStudy (anonymous.):

y+2 = 3x -6 y=3x-7 ?

OpenStudy (anonymous.):

Y=3+7 :D !!

OpenStudy (anonymous):

so it's A) :)

OpenStudy (anonymous):

congrats

OpenStudy (anonymous.):

Thank you !!